OpenStudy (anonymous):

What is the volume of a cone with a height of 6 centimeters and a radius of 2 centimeters (use 3.14 for )? 25.12 cubic centimeters 24.12 cubic centimeters 25.24 cubic centimeters 24.00 cubic centimeters

OpenStudy (amistre64):

cone volumes are: (1/3) (base area) (height)

OpenStudy (anonymous):

volume of a cone = 1/3 volume of the cylinder. so (3.14)(2^2)(6)(1/3) = 25.12

OpenStudy (amistre64):

6/3 = 2 right off the bat.. 2 (pi 2^2) = 8 pi